OpenStudy (ashley1nonly):

x^2+y^2+z^2+3x-4z+1=0 (x^2+3x)+y^2+(z^2-4z)=-1 where do I go from here to get the radius and center for the sphere (x-x0)^2+(y-y0)^2+(z-z0)^2=a^2

OpenStudy (irishboy123):

from \(x^2+y^2+z^2+3x-4z+1=0\) to \((x^2+3x)+y^2+(z^2-4z)=-1\) you need to complete the squares for x,y,z so, for x: \( x^2+3x= (x + \frac{3}{2})^2 - \frac{9}{4} \) then add them all up :p

OpenStudy (ashley1nonly):

where did 3/2 come from.

OpenStudy (irishboy123):

what is \( (x + \frac{3}{2})^2\)??

OpenStudy (ashley1nonly):

oh okay thanks